# Optimization via Quantum Preconditioning

### v1

@@ I Introduction
[…]
@@
The advent of quantum technologies as a novel computing paradigm [13, 14] has sparked interest in exploring whether discrete optimization can leverage quantum phenomena for improved efficiency [15]. On the one hand, quantum adiabaticity [16, 17] serves as the foundation for quantum annealers and, more broadly, analog quantum computation [18, 19]. Analog platforms, through empirical observation, have demonstrated an asymptotic scaling advantage over classical methods such as SA and parallel tempering on specific classes of problems [20, 21, 22, 23]. However, an absolute advantage in computational time has not yet been realized at currently addressed problem sizes on current quantum hardware. On the other hand, digital quantum computers operating via programmable quantum logical operations typically leverage the variational principle to address discrete optimization [24]. In that context, a relevant algorithm is the quantum approximate optimization algorithm (QAOA) [25, 26]. In certain limiting cases and for specific problem types, the QAOA can provide performance guarantees on the accuracy of the returned solution [25, 27, 28, 29, 30]. Additionally, the QAOA can be used as a heuristic, where classical emulations support evidence of a scaling advantage for certain problems that is yet to be realized in practice [31, 32]. Besides the QAOA, other quantum algorithms with a potential scaling advantage have been advanced [33, 34, 35, 36]. While the evidence for a scaling advantage is promising, transforming it into a tangible advantage for quantum computers to deliver utility remains a substantial engineering challenge. Alongside hardware development, further progress in quantum algorithms is equally important in this endeavor.
@@
Here, we introduce the idea of quantum preconditioning for discrete optimization problems (Section II). In the classical context, preconditioning seeks to transform a problem into a more suitable form for a solver. It aims to make the preconditioned problem easier to solve than its original version by, e.g., converging faster to a solution. A notable example arises in state-of-the-art classical linear system solvers, where classical preconditioning is instrumental to their performance [37, 38]. A related idea for solving mixed integer programming problems is “presolving” [39, 40]: It encompasses preprocessing techniques applied to the input problem to make it more amenable for a solver.
@@
We propose a quantum preconditioner for quadratic unconstrained binary optimization problems [41, 42] based on the QAOA with $p$ layers which works as follows. The goal is to extremize a quadratic objective function $C(\boldsymbol{z})$ over $N$ discrete variables $\boldsymbol{z}=(z_{1},z_{2},\ldots,z_{N})^{T}$, where the objective function is encoded by an adjacency matrix $\mathsf{W}\in\mathbb{R}^{N\times N}$. First, we execute the QAOA on the original $N$-variable problem encoded by $\mathsf{W}$. Then, we use the resulting quantum state to estimate two-point correlations between qubits as a pairwise correlation matrix $\mathsf{Z}^{(p)}\in\mathbb{R}^{N\times N}$. Finally, we substitute the adjacency matrix $\mathsf{W}$ by the correlation matrix $\mathsf{Z}^{(p)}$ in the problem’s objective function. Hence, the quantum preconditioner preserves the structure of quadratic unconstrained binary optimization and the correlation matrix is the quantum-preconditioned problem solved in place of the original problem by a suitable classical solver. This allows classical solvers to work with either the original input or a preconditioned version, potentially leading to a direct improvement over the current state-of-the-art classical approach.
@@
Building on the adiabatic theorem, we demonstrate that in the adiabatic limit ($p\to+\infty$[16], the quantum transformation renders solving the problem optimally a trivial task. Any local updates that improve a candidate solution $\boldsymbol{z}$ lead to the optimal solution $\boldsymbol{z}_{\textbf{opt}}$, thereby benefiting classical solvers like SA and parallel tempering. In the opposite limit, at a shallow circuit depth $p\leq 2$, we perform classical emulations of the QAOA to test the performance of proposed quantum preconditioner. We find that best-in-class classical heuristics can converge more rapidly on an average case. In particular, SA and the BM algorithm can benefit from low-depth quantum preconditioning in solving a range of problems, such as Sherrington-Kirkpatrick spin glasses [43], random $3$-regular graph maximum-cut problems, and a real-world grid energy problem [44].
@@
Using a light-cone decomposition technique [45], we benchmark the proposed approach on problems with thousands of variables and investigate the effect of increased circuit depth on its convergence. For instance, on a ensemble of random $3$-regular graph maximum-cut problems with $N=4,096$ variables, quantum preconditioning can make BM and SA converge one order of magnitude faster to a typical solution $\boldsymbol{z}$ with an average approximation ratio $\alpha=C(\boldsymbol{z})/C(\boldsymbol{z}_{\textbf{opt}})=99.9\%$ (Section III). When accounting for the additional time taken for preconditioning, the benefit offered by shallow circuits remains and translates into a practical advantage for random $3$-regular graph maximum-cut problems through classical QAOA emulations. To the best of our knowledge, this is the first reported quantum-inspired advantage for optimization. The improved convergence observed upon increasing the circuit depth opens the door to a hardware-based quantum advantage using circuits that are beyond the reach of classical emulators for preconditioning.
@@
In addition, we study why quantum preconditioning makes the problem easier by considering small-$N$ problems at large $p$ (Appendix B). We test an experimental implementation of quantum preconditioning of the real-world grid energy problem [44] on Rigetti Ankaa-3 superconducting device (Appendix C). Finally, We identify challenges and discuss prospects for quantum preconditioning to deliver utility (Section IV). We give our conclusions in Section V.
@@ II Quantum Preconditioning
We consider discrete optimization problems of the form [41, 42]
@@
where $\hat{H}$, $\hat{X}$, and $\langle\boldsymbol{z}|\hat{C}|\boldsymbol{z^{\prime}}\rangle=C(\boldsymbol{z})\delta_{\boldsymbol{z}\boldsymbol{z^{\prime}}}$ are the Hadamard gate, Pauli-$X$ operator, and operator encoding the objective function, respectively. There is a one-to-one mapping between a variable and a qubit. Eq. (3) corresponds to a standard $p$-layer QAOA circuit [25, 26] parameterized via $\gamma_{\ell}$ and $\beta_{\ell}$, which are optimal when extremizing $\langle\hat{C}\rangle_{p}$. In a similar form to the original problem of Eq. (1), the preconditioned problem’s objective function reads $\tilde{C}^{(p)}(\boldsymbol{z})=\frac{1}{2}\boldsymbol{z}^{T}\mathsf{Z}^{(p)}\boldsymbol{z}$.
@@ II.1 The Asymptotic Infinite-Circuit-Depth Limit
At infinite depth ($p\to+\infty$), the QAOA converges to the optimal solution [25]. In the presence of a doubly degenerate optimal solution ±$\pm\boldsymbol{z}_{\textbf{opt}}$ due to the $\mathbb{Z}_{2}$ global sign-flip symmetry in the objective function of Eq. (1), the preconditioner takes the form $\mathsf{Z}^{(\infty)}=\mathsf{I}-\boldsymbol{z}_{\textbf{opt}}\boldsymbol{z}_{\textbf{opt}}^{T}$, where $\mathsf{I}$ is an $N\times N$ identity matrix. The objective function of the quantum preconditioned problem now reads
@@
which is minimized for $\boldsymbol{z}=\pm\boldsymbol{z}_{\textbf{opt}}$. Therefore, both the original and preconditioned problems in the infinite-depth limit share the same optimal solution. However, while $\boldsymbol{z}_{\textbf{opt}}$ is nontrivial to find based on the original problem of Eq. (1) in the general case, it can be straightforwardly obtained via local steps based on Eq. (4). Indeed, $\tilde{C}^{(\infty)}(\boldsymbol{z}_{\textbf{opt}})=-N(N-1)$ marks the absence of frustration in the quantum preconditioned problem. Each variable $z_{i}$ can be set to ±$\pm 1$ to minimize locally each term $\mathsf{Z}^{(\infty)}_{ij}z_{i}z_{j}$, which minimizes the objective function globally. This signals the absence of glassy properties [46, 47, 48] in the quantum preconditioned problem in the infinite-depth limit. Hence, we expect a strong benefit for solvers, such as SA [6], parallel tempering [7], and others, that rely on local steps for convergence. This is evidenced by the decrease in frustration index of the preconditioned problem with QAOA circuit depth $p$, as shown in Appendix B for small-$N$ problem instances.
To extend the infinite-depth properties of quantum preconditioning to problems beyond doubly degenerate optimal solution ±$\pm\boldsymbol{z}_{\textbf{opt}}$, another preconditioner is needed. For instance, for problems with a unique solution $\boldsymbol{z}_{\textbf{opt}}$ due to an extra nonzero term of the form $\sum_{i}h_{i}z_{i}$ with $h_{i}\in\mathbb{R}$, one may use $\sum_{i}h_{i}z_{i}\leftarrow-\sum_{i}\langle\hat{Z}_{i}\rangle_{p}z_{i}$. For problems with degenerate optimal solutions beyond ±$\pm\boldsymbol{z}_{\textbf{opt}}$, one can add vanishing one-body terms $\lim_{h_{i}\to 0}h_{i}z_{i}$ in the original objective function to favor a single solution.
Figure 1: Average performance for the maximum-cut problem on $N$-variable random $3$-regular graphs via the classical simulated annealing (SA) solver based on the original and quantum-preconditioned problems. (A) Workflow diagram of quantum preconditioning. (B) Average run-time for SA to get to an approximation ratio of $99.9\%$ as a function of $N$. (C) Average approximation ratio versus SA run-time for $N=4,096$. Each data point is averaged over $200$ randomly generated problem instances. Run-times corresponds to a 64Gb MacBook Pro with an Apple M1 Max chip. Error bars indicate the standard error of the mean.
@@ II.2 At a Shallow Circuit Depth
The performance at practical circuit depths may be independent of these infinite-depth considerations. In the following, we investigate the performance of the preconditioner of Eq. (2) on problems with at least a doubly degenerate optimal solution ±$\pm\boldsymbol{z}_{\textbf{opt}}$ (potentially more) due to their global $\mathbb{Z}_{2}$ sign-flip symmetry. Here, the optimal solution for $\mathsf{Z}$ is generically different from the optimal solution for $\mathsf{W}$. Nevertheless, we find that this form of quantum preconditioning at shallow depth can still provide a benefit. Quantifying the quality of a solution $\boldsymbol{z}$ through the approximation ratio $\alpha=C(\boldsymbol{z})/C(\boldsymbol{z}_{\textbf{opt}})$, we find that the classical solvers achieve $\alpha\approx 1$ when working with the preconditioned problem, but often do so with fewer iterations than with the original problem, which may translate to lesser computational time. This is the central result of this work. We emphasize that, although a solution $\boldsymbol{z}$ is obtained by solving the original or preconditioned problem, the objective function $C(\boldsymbol{z})$ and the optimal solution $\boldsymbol{z}_{\textbf{opt}}$ are always with respect to the original problem for evaluating the approximation ratio $\alpha$.
The intuition behind the QAOA preconditioner of Eq. (2) at a shallow circuit depth $p$ is that the correlation matrix contains similar information to the original problem. Indeed, locally, two variables $z_{i}$ will tend to take values such that $\mathsf{W}_{ij}z_{i}z_{j}$ is minimized, which translates into $\textrm{sign}(\mathsf{W}_{ij})\simeq-\textrm{sign}\langle\hat{Z}_{i}\hat{Z}_{j}\rangle$. Thus, one may anticipate that solving a problem based on $\mathsf{W}_{ij}$ or $\mathsf{Z}_{ij}^{(p)}$ leads to similar solutions. A special case of quantum preconditioning is the quantum relax-and-round algorithm described in Refs. [49, 45]. There, it was a shown that a classical relax-and-round solver yields identical solutions for the original problem given by $\mathsf{W}$ and the preconditioned problem at $p=1$ given by $\mathsf{Z}^{(p=1)}$ on a range of problems, including Sherrington-Kirkpatrick spin glasses and random $3$-regular graph maximum-cut problems.
The same intuition holds true for any classical solver. Here, we employ the state-of-the-art SA and BM solvers to solve Sherrington-Kirkpatrick spin glasses [43], random $3$-regular graph maximum-cut problems, and a real-world grid energy problem [44]. We find that at small $p$, these solvers converge faster to a given value of the approximation ratio $\alpha$ on the quantum-preconditioned versus the original problems, and conversely, yield a higher approximation ratio $\alpha$ in a given amount of time. In addition, we investigate in Appendix B large-$p$ quantum preconditioning on small-$N$ problem instances.

### v2

@@ I Introduction
Optimization problems over discrete variables are ubiquitous [1, 2, 3]; however, they continue to pose a significant challenge for traditional computing paradigms. The absence of provably efficient algorithms in the general case unless $\textrm{P}=\textrm{NP}$ [4, 5] makes heuristic solvers the practical method of choice. Celebrated examples include physics-inspired approaches, such as simulated annealing (SA) [6] or parallel tempering [7], as well as heuristics that build on approximate methods with performance guarantees, like the Burer-Monteiro (BM) solver [8, 9] based on the Goemans-Williamson (GW) algorithm for the maximum-cut problem [10, 11]. These heuristics represent the state of the art in discrete optimization on a range of problems [12].
@@
The advent of quantum technologies as a new computing paradigm [13, 14] has sparked interest in exploring whether discrete optimization can leverage quantum phenomena for improved efficiency [15]. On the one hand, quantum adiabaticity [16, 17] serves as the foundation for quantum annealers and, more broadly, analog quantum computation [18, 19]. Analog platforms, through empirical observation, have demonstrated an asymptotic scaling advantage over classical methods such as SA and parallel tempering on specific classes of problems [20, 21, 22, 23]; however, an absolute advantage in computational time has not yet been realized at currently addressed problem sizes on current quantum hardware. On the other hand, digital quantum computers operating via programmable quantum logical operations typically leverage the variational principle to address discrete optimization [24]. In that context, a relevant algorithm is the quantum approximate optimization algorithm (QAOA) [25, 26]. In certain limiting cases and for specific problem types, the QAOA can provide performance guarantees on the accuracy of the returned solution [25, 27, 28, 29, 30]. Additionally, the QAOA can be used as a heuristic, where classical emulations support evidence of a scaling advantage for certain problems that is yet to be realized in practice [31, 32]. In addition to the QAOA, other quantum algorithms with a potential scaling advantage have been advanced [33, 34, 35, 36]. While the evidence for a scaling advantage is promising, transforming this into a tangible advantage for quantum computers to deliver utility remains a substantial engineering challenge. Alongside hardware development, further progress in quantum algorithms is equally important in this endeavor.
@@
Here, we introduce the idea of quantum preconditioning for discrete optimization problems (Sec. II). In the classical context, preconditioning seeks to transform a problem into a more suitable form for a solver. It aims to make the preconditioned problem easier to solve than its original version by, e.g., converging faster to a solution. A notable example arises in state-of-the-art classical linear system solvers, where classical preconditioning is instrumental to their performance [37, 38]. A related idea for solving mixed integer programming problems is “presolving” [39, 40]: this involves preprocessing techniques being applied to the input problem to make it more tractable for a solver.
@@
We propose a quantum preconditioner for quadratic unconstrained binary optimization problems [41] based on the QAOA with $p$ layers which works as follows. The goal is to extremize a quadratic objective function $C(\boldsymbol{z})$ over $N$ discrete variables $\boldsymbol{z}=(z_{1},z_{2},\ldots,z_{N})^{T}$, where the objective function is encoded by an adjacency matrix $\mathsf{W}\in\mathbb{R}^{N\times N}$. First, we execute the QAOA on the original $N$-variable problem encoded by $\mathsf{W}$. Then, we use the resulting quantum state to estimate two-point correlations between qubits as a pairwise correlation matrix $\mathsf{Z}^{(p)}\in\mathbb{R}^{N\times N}$. Finally, we replace the adjacency matrix $\mathsf{W}$ with the correlation matrix $\mathsf{Z}^{(p)}$ in the problem’s objective function. Hence, the quantum preconditioner preserves the structure of quadratic unconstrained binary optimization, and the correlation matrix is the quantum-preconditioned problem solved in place of the original problem using a suitable classical solver. This allows classical solvers to work with either the original input or a preconditioned version, potentially leading to a direct improvement over the current state-of-the-art classical approach.
@@
Building on the adiabatic theorem, we demonstrate that in the adiabatic limit ($p\to+\infty$[16], the quantum transformation renders solving the problem optimally a trivial task. Any local updates that improve a candidate solution $\boldsymbol{z}$ lead to the optimal solution $\boldsymbol{z}_{\textbf{opt}}$, thereby benefiting classical solvers like SA and parallel tempering. In the opposite limit, at a shallow circuit depth $p\leq 2$, we perform classical emulations of the QAOA to test the performance of the proposed quantum preconditioner. We find that best-in-class classical heuristics can converge more rapidly on an average case. In particular, SA and the BM algorithm can benefit from low-depth quantum preconditioning in solving a range of problems, such as Sherrington-Kirkpatrick spin glasses [42], random $3$-regular graph maximum-cut problems, and a real-world grid energy problem [43].
@@
Using a light-cone decomposition technique [44], we benchmark the proposed approach on problems with thousands of variables and investigate the effect of increased circuit depth on its convergence. For instance, on an ensemble of random $3$-regular graph maximum-cut problems with $N=4,096$ variables, quantum preconditioning can make BM and SA converge one order of magnitude faster to a typical solution $\boldsymbol{z}$ with an average approximation ratio $\alpha=C(\boldsymbol{z})/C(\boldsymbol{z}_{\textbf{opt}})=99.9\%$ (Sec. III). When accounting for the additional time taken for preconditioning, the benefit offered by shallow circuits remains and translates into a practical advantage for random $3$-regular graph maximum-cut problems through classical QAOA emulations. To the best of our knowledge, this is the first reported quantum-inspired advantage for optimization. The improved convergence observed upon increasing the circuit depth opens the door to a hardware-based quantum advantage using circuits that are beyond the reach of classical emulators for preconditioning.
@@
In addition, we study why quantum preconditioning makes the problem easier by considering small-$N$ problems at large $p$ (Appendix B). We test an experimental implementation of quantum preconditioning of the real-world grid energy problem [43] on a Rigetti Ankaa-3 superconducting device (Appendix D). Finally, we identify challenges and discuss prospects for quantum preconditioning to deliver utility (Sec. IV). We give our conclusions in Sec. V.
@@ II Quantum Preconditioning
We consider discrete optimization problems of the form [41]
@@
where $\hat{H}$, $\hat{X}$, and $\langle\boldsymbol{z}|\hat{C}|\boldsymbol{z^{\prime}}\rangle=C(\boldsymbol{z})\delta_{\boldsymbol{z}\boldsymbol{z^{\prime}}}$ are the Hadamard gate, Pauli-$X$ operator, and operator encoding the objective function, respectively. There is a one-to-one mapping between a variable and a qubit. Equation (3) corresponds to a standard $p$-layer QAOA circuit [25, 26] parameterized via $\gamma_{\ell}$ and $\beta_{\ell}$, which are optimal when extremizing $\langle\hat{C}\rangle_{p}$. In a similar form to the original problem of Eq. (1), the preconditioned problem’s objective function reads $\tilde{C}^{(p)}(\boldsymbol{z})=\frac{1}{2}\boldsymbol{z}^{T}\mathsf{Z}^{(p)}\boldsymbol{z}$. In the following, quantum preconditioning is performed at optimal or near-optimal $\gamma_{\ell}$ and $\beta_{\ell}$ values with respect to extremizing $\langle\hat{C}\rangle_{p}$ via the QAOA. Hence, the computational cost of preconditioning is the same as executing the QAOA. No additional information than what is available as part of the standard QAOA is needed.
@@ II.1 The asymptotic infinite-circuit-depth limit
At infinite depth ($p\to+\infty$), the QAOA converges to the optimal solution [25]. In the presence of a doubly degenerate optimal solution $\pm\boldsymbol{z}_{\textbf{opt}}$ due to the $\mathbb{Z}_{2}$ global sign-flip symmetry in the objective function of Eq. (1), the preconditioner takes the form $\mathsf{Z}^{(\infty)}=\mathsf{I}-\boldsymbol{z}_{\textbf{opt}}\boldsymbol{z}_{\textbf{opt}}^{T}$, where $\mathsf{I}$ is an $N\times N$ identity matrix. The objective function of the quantum-preconditioned problem now reads
@@
which is minimized for $\boldsymbol{z}=\pm\boldsymbol{z}_{\textbf{opt}}$. Therefore, both the original and preconditioned problems in the infinite-depth limit share the same optimal solution; however, while $\boldsymbol{z}_{\textbf{opt}}$ is nontrivial to find based on the original problem of Eq. (1) in the general case, it can be straightforwardly obtained via local steps based on Eq. (4). Indeed, $\tilde{C}^{(\infty)}(\boldsymbol{z}_{\textbf{opt}})=-N(N-1)$ marks the absence of frustration in the quantum-preconditioned problem. Each variable $z_{i}$ can be set to $\pm 1$ to locally minimize each term $\mathsf{Z}^{(\infty)}_{ij}z_{i}z_{j}$, which minimizes the objective function globally. This signals the absence of glassy properties [45, 46, 47] in the quantum-preconditioned problem in the infinite-depth limit. Hence, we expect a strong benefit for solvers, such as SA [6], parallel tempering [7], and others, that rely on local steps for convergence. This is evidenced by the decrease in frustration index of the preconditioned problem with QAOA circuit depth $p$, as shown in Appendix B for small-$N$ problem instances.
To extend the infinite-depth properties of quantum preconditioning to problems beyond the doubly degenerate optimal solution $\pm\boldsymbol{z}_{\textbf{opt}}$, another preconditioner is needed. For instance, for problems with a unique solution $\boldsymbol{z}_{\textbf{opt}}$ due to an extra nonzero term of the form $\sum_{i}h_{i}z_{i}$ with $h_{i}\in\mathbb{R}$, one may use $\sum_{i}h_{i}z_{i}\leftarrow-\sum_{i}\langle\hat{Z}_{i}\rangle_{p}z_{i}$. For problems with degenerate optimal solutions beyond $\pm\boldsymbol{z}_{\textbf{opt}}$, one can add vanishing one-body terms $\lim_{h_{i}\to 0}h_{i}z_{i}$ in the original objective function to favor a single solution.
Figure 1: Average performance for the maximum-cut problem on $N$-variable random $3$-regular graphs via the classical simulated annealing (SA) solver based on the original and quantum-preconditioned problems. (A) Workflow diagram of quantum preconditioning. (B) Average run-time for SA to get to an approximation ratio of $99.9\%$ as a function of $N$. (C) Average approximation ratio versus SA run-time for $N=4,096$. Each data point is averaged over $200$ randomly generated problem instances. Run-times correspond to a 64GB MacBook Pro with an Apple M1 Max chip. Error bars indicate the standard error of the mean.
@@ II.2 At a shallow circuit depth
The performance at practical circuit depths may be independent of these infinite-depth considerations. In the following, we investigate the performance of the preconditioner of Eq. (2) on problems with at least a doubly degenerate optimal solution $\pm\boldsymbol{z}_{\textbf{opt}}$ (potentially more) due to their global $\mathbb{Z}_{2}$ sign-flip symmetry. Here, the optimal solution for $\mathsf{Z}$ is generically different from the optimal solution for $\mathsf{W}$. Nevertheless, we find that this form of quantum preconditioning at shallow depth can still provide a benefit. Quantifying the quality of a solution $\boldsymbol{z}$ through the approximation ratio $\alpha=C(\boldsymbol{z})/C(\boldsymbol{z}_{\textbf{opt}})$, we find that the classical solvers achieve $\alpha\approx 1$ when working with the preconditioned problem, but they often do so with fewer iterations than with the original problem, which may translate to less computational time. This is the central result of this work. We emphasize that, although a solution $\boldsymbol{z}$ is obtained by solving the original or preconditioned problem, the objective function $C(\boldsymbol{z})$ and the optimal solution $\boldsymbol{z}_{\textbf{opt}}$ are always with respect to the original problem for evaluating the approximation ratio $\alpha$.
The intuition behind the QAOA preconditioner of Eq. (2) at a shallow circuit depth $p$ is that the correlation matrix contains similar information to the original problem. Indeed, locally, two variables $z_{i}$ will tend to take values such that $\mathsf{W}_{ij}z_{i}z_{j}$ is minimized, which translates into $\textrm{sign}(\mathsf{W}_{ij})\simeq-\textrm{sign}\langle\hat{Z}_{i}\hat{Z}_{j}\rangle$. Thus, one may anticipate that solving a problem based on $\mathsf{W}_{ij}$ or $\mathsf{Z}_{ij}^{(p)}$ will lead to similar solutions. A special case of quantum preconditioning is the quantum relax-and-round algorithm described in Refs. [48, 44]. There, it was a shown that a classical relax-and-round solver yields identical solutions for the original problem given by $\mathsf{W}$ and the preconditioned problem at $p=1$ given by $\mathsf{Z}^{(p=1)}$ on a range of problems, including Sherrington-Kirkpatrick spin glasses and random $3$-regular graph maximum-cut problems.
The same intuition holds true for any classical solver. Here, we employ the state-of-the-art SA and BM solvers to solve Sherrington-Kirkpatrick spin glasses [42], random $3$-regular graph maximum-cut problems, and a real-world grid energy problem [43]. We find that at small $p$, these solvers converge faster to a given value of the approximation ratio $\alpha$ on the quantum-preconditioned versus the original problems, and conversely, they yield a higher approximation ratio $\alpha$ in a given amount of time. In addition, in Appendix B, we investigate large-$p$ quantum preconditioning on small-$N$ problem instances.